OpenStudy (anonymous):

How do you factor x^2-4x+24

OpenStudy (mertsj):

It won't factor by the usual methods since there are not 2 numbers whose product is 24 and whose sum is -4

OpenStudy (anonymous):

can't be factored. use the quadratic formula which is x= -b + or - the square root of b^2 - 4ac all over 2a a= the x^2 which is 1 b=the -4 c=24 hope that helped :)

OpenStudy (anonymous):

I don't know what to do, though.

OpenStudy (mertsj):

So if you want to factor it, you could use the quadratic formula and find the roots. Then using the roots, you could write the factors.

OpenStudy (anonymous):

I think that helped

OpenStudy (mertsj):

Are you sure you copied it right?

OpenStudy (anonymous):

Yes

OpenStudy (mertsj):

\[x=\frac{4\pm \sqrt{16-4(24)}}{2}=\frac{4\pm \sqrt{-80}}{2}=\frac{4\pm4\sqrt{5}i}{2}=2\pm2\sqrt{5}i\]

OpenStudy (mertsj):

So the factors must by: \[[x-(2+2\sqrt{5}i)][x-(2-2\sqrt{5}i)]\]